# Is This Collection Worth My LLM’s Time? Automatically Measuring Information Potential in Text Corpora

### v1

@@ Abstract
As large language models (LLMs) converge towards similar capabilities, the key to advancing their performance lies in identifying and incorporating valuable new information sources. However, evaluating which text collections are worth the substantial investment required for digitization, preprocessing, and integration into LLM systems remains a significant challenge. We present a novel approach to this challenge: an automated pipeline that evaluates the potential information gain from text collections without requiring model training or fine-tuning. Our method generates multiple choice questions (MCQs) from texts and measures an LLM’s performance both with and without access to the source material. The performance gap between these conditions serves as a proxy for the collection’s information potential. We validate our approach using three strategically selected datasets: EPFL PhD manuscripts (likely containing novel specialized knowledge), Wikipedia articles (presumably part of training data), and a synthetic baseline dataset. Our results demonstrate that this method effectively identifies collections containing valuable novel information, providing a practical tool for prioritizing data acquisition and integration efforts.
@@ 1 Introduction
    A self-supervised methodology for ensuring MCQ quality through complementary similarity metrics that address two key challenges: verifying question relevance and grounding in source material, while ensuring distractors are plausible yet unambiguously incorrect
@@ 2 Related Work
Information Gain. The challenge of efficiently selecting new information sources for LLMs can be viewed through the lens of optimal experiment design, a field pioneered by Fedorov et al. (1972). This framework emphasizes maximizing information gain while being strategic about resource allocation – a particularly relevant consideration given the costs associated with integrating new data into LLM systems. Information gain itself has been conceptualized across various fields: in information theory, it relates to reductions in algorithmic information content Cover et al. (1989); in machine learning, it quantifies a feature’s contribution to model performance Odhiambo Omuya et al. (2021); and in cognitive science, it represents uncertainty reduction in our experience of the world Damiano et al. (2021). While these theoretical frameworks provide valuable insights, they have nott been previously applied to the specific challenge of evaluating the potential value of text collections for enhancing LLM knowledge. Our work bridges this gap by proposing a practical, MCQ-based approach that quantifies information gain by measuring an LLM’s ability to answer questions about a text collection with and without access to the source material.
@@ 3 Methods
Generating high-quality MCQs presents unique challenges, particularly due to the absence of ground truth signals typically available in human-curated educational assessments. Without explicit supervision on what constitutes a good question or appropriate distractors, we must derive reliable metrics to filter and validate the generated MCQs. This self-supervised setting raises two critical challenges: ensuring questions are both relevant and non-trivial, and guaranteeing that distractors are plausible yet unambiguously incorrect. The latter is particularly crucial when evaluating LLMs – if distractors are too simple, models might succeed through elimination rather than true knowledge; if too similar to the correct answer, they might confuse models even when provided with context. To address these challenges, we propose similarity-based filtering metrics that serve as proxy signals for question quality and difficulty, enabling automated quality control without human supervision.
@@ 3.3 Evaluation
Here, $\mathbbm{1}(\cdot)$ is an indicator function equal to 1 if the condition is true, and 0 otherwise. The denominator ensures that we focus on questions t rather than letting widespread model failures distort the results.
Position Bias Mitigation. To address the known issue of positional bias in LLMs Shi et al. (2024a), we evaluate each MCQ four times, rotating the correct answer through all possible positions (A, B, C, D) while randomly arranging the distractors. This effect can be seen in Figures 4, 6, and 5. This ensures that model performance reflects true knowledge rather than position-based preferences.
@@ 3.4 Data
To evaluate our pipeline’s efficacy, we conduct experiments across three strategically selected datasets representing distinct knowledge domains: (1) 177 EPFL PhD manuscripts containing specialized academic research, (2) Wikipedia articles, and (3) synthetic LLM-generated text. This dataset composition enables systematic assessment of knowledge novelty. We hypothesize that PhD manuscripts will yield the highest knowledge gaps, containing recent research contributions likely absent from training data. The synthetic dataset, comprising model-generated content, serves as a lower-bound baseline for knowledge novelty. Wikipedia articles, being a primary source for LLM training Brown et al. (2020); Touvron et al. (2023), function as a control group representing knowledge presumably well-embedded in the models’ parameters.
Data collection. The Wikipedia dataset is created by leveraging the Wikipedia API. With the help of the PhD manuscript titles, related articles are fetched and combined into similar-sized texts. Similarly, the manuscript titles are used to generate the baseline dataset. To overcome the LLMs’ generation limits, in a first step, subtopics are generated surrounding the manuscript titles. Then, chunks of around 600 words are generated, using the subtopics as overarching themes, and concatenated.
@@ 4 Results
Our analysis reveals three key findings: (1) similarity thresholding effectively increases question difficulty while maintaining answerable questions, (2) PhD manuscripts contain significantly more novel information compared to Wikipedia and synthetic datasets, and (3) larger language models show consistently lower information potential across all datasets, suggesting better knowledge retention during pre-training. We present detailed evidence for each of these findings below.
@@ 4.1 Effectiveness of Similarity Thresholding
(a) Only Thresholding with Cosine Similarity
(b) Only Thresholding with Jaccard and Rouge-L
Figure 2: Llama 3 70B Performance on EPFL Dataset Across Different Cutoff Percentages. (with Separated Thresholding and Number of MCQs for the Corresponding Percentiles). NC 4x: Evaluation with no context correct for all 4 bias mitigation evaluations of each question; and WC 4x: Evaluation with context correct for all 4 bias mitigation evaluations of each question.
(a) Llama 70B Baseline
IP: 0.125
(b) Llama 70B Wikipedia
IP: 0.136
(c) Llama 70B EPFL
IP: 0.229
(d) GPT-4o Wikipedia
IP: 0.110
(e) GPT-4o EPFL
IP: 0.211
(f) Legend
Figure 3: Information Potential Analysis Across Datasets and Models. Venn diagrams showing correct response overlap between context-free and context-provided conditions. Higher Information Potential (IP) scores indicate greater novel information content, with PhD manuscripts (EPFL) showing consistently higher IP (0.211-0.229) compared to Wikipedia (0.110-0.136) and synthetic baseline (0.125). Both open and closed-source models exhibit similar patterns despite architectural differences.
Effect of Cosine Similarity Thresholding. Figure 2(a) displays the effect of cosine thresholding. It shows a strong decrease of around 10% between the absence of thresholding and the 50th percentile cutoff in the performance of the model with no context (NC 4x), while that of the model with context (WC 4x) remains steady. Therefore, the cosine similarity has the expected effect of increasing the MCQ difficulty when answering without context while maintaining the context-based performance.
Effect of Jaccard and ROUGE-L Thresholding. A different trend is observed when applying the Jaccard and ROUGE-L thresholding, as seen in Figure 2(b). Here, a slight increase in performance with context (WC 4x) of 2% can be observed between no thresholding and the 50th percentile cutoff, while performance without context (NC 4x) remains stable.
Comparative Analysis of Thresholding Methods. Our results demonstrate that cosine similarity thresholding is more effective at increasing question difficulty than Jaccard and ROUGE-L methods. This difference likely stems from cosine similarity’s ability to capture semantic relationships in the embedded space, while Jaccard and ROUGE-L focus on surface-level textual overlap. The gradual reduction in MCQ count under cosine thresholding (from 1633 to 653) suggests a more controlled filtering process compared to the sharp drop-off observed with Jaccard and ROUGE-L (907 to 479 MCQs between 50th and 60th percentiles).
@@ 4.2 Information Potential (IP) Across Different Datasets
EPFL PhD Manuscripts (IP: 0.229). The EPFL dataset yields the highest information potential among all tested collections. We observe a significant performance gap between context-free (73.4%) and context-provided (98%) conditions, highlighting the substantial amount of novel information in these documents. The questions generated from this dataset predominantly focus on specialized technical concepts and novel research findings. These performance patterns strongly indicate that PhD manuscripts contain substantial novel information not present in the model’s pre-training data.
Wikipedia Dataset (IP: 0.136). The Wikipedia dataset demonstrates moderate information potential, with a notably smaller gap between context-free and context-provided performance (83.6% vs. 97.8%). Questions from this dataset primarily cover general knowledge and well-documented facts. The performance patterns suggest partial inclusion of this content in the model’s training data, which aligns with our expectations given that these articles were selected from versions predating the model’s training cutoff date.
@@ 4.4 Qualitative Analysis of High-Value Information
Table 1: Examples of technical terminology, novel findings, and complex relationships in EPFL PhD manuscripts. MCQs requiring context for correct model responses
@@
    Novel Findings: The question regarding "interference errors" exemplifies how PhD manuscripts capture recent research outcomes. The detailed distinction between interference errors and reduction of intentionality errors represents novel academic insights that weren’t available during model pre-training. This type of question effectively identifies valuable new knowledge contributions from academic manuscripts. The consistency with which models fail these questions without context, despite their strong general reasoning capabilities, suggests genuine knowledge gaps rather than reasoning limitations.
@@ 5.1 Summary of Contributions
    Empirical Validation: Our results validate the method’s effectiveness by demonstrating alignment with intuitive expectations across both dataset types and model scales. The information potential increases from synthetic baseline (0.125) to Wikipedia (0.136) to EPFL manuscripts (0.229), while larger models (GPT-4o) consistently show lower information potential than smaller ones (Llama 70B), suggesting better knowledge retention.
@@
5.2 Limitations and Future Work
Practical Considerations. Our approach, while effective, presents several methodological limitations that suggest directions for future research. The current pipeline’s generation-first approach, where questions are created before applying filtering criteria, could be enhanced by incorporating context-answer alignment and distractor plausibility directly into the generation process. Such an integrated approach could improve question quality while reducing computational overhead.
The scope of our evaluation faces an inherent challenge that manifested both as a limitation of our study and, ironically, in the process of writing this paper: identifying datasets that are definitively outside of LLMs’ training data. While PhD manuscripts serve as a reasonable proxy for novel information, the rapid integration of online content into training sets makes finding suitable test collections increasingly difficult. This challenge underscores the timeliness of our work, as it becomes crucial to identify and preserve valuable information sources before they are absorbed into general training data.
Perhaps most significantly, our method deliberately focuses on information selection rather than distillation. This choice, while aligned with our goal of evaluating potential value, leaves open questions about how selected information can be optimally incorporated into LLM training. Recent work has begun to address this gap, with approaches such as SKILL (Moiseev et al., 2022) exploring efficient knowledge integration methods. However, Liu et al. (2024) highlight potential conflicts between newly added information and existing knowledge compressed in LLMs’ internal representations, suggesting that information integration remains a complex challenge requiring further investigation.
@@ 5.2 Limitations and Future Work
Our effort to categorize high value information sheds light on fundamental challenges in distinguishing between information that is unknown to an LLM versus information it fails to retrieve. This raises important questions about the relationship between knowledge possession and practical competence in AI systems. When an LLM consistently fails to demonstrate knowledge in a specific domain, the distinction between these cases may become less relevant from a practical perspective. This observation has broader implications for how we conceptualize and evaluate knowledge in AI systems, suggesting that performance-based metrics might be more meaningful than theoretical attempts to map internal knowledge representations. Furthermore, this highlights the need for a more nuanced understanding of how LLMs process and utilize information, particularly when developing strategies for knowledge integration and model enhancement.
@@ A.2.1 Positional Bias
(a) Baseline
(b) Wikipedia
(c) Venice
@@
(a) EPFL - No Context
(b) EPFL - With Context
(c) Wikipedia - No Context
(d) Wikipedia - With Context
@@
(a) EPFL - No Context
(b) EPFL - With Context
(c) Wikipedia - No Context
(d) Wikipedia - With Context
(e) Baseline - No Context
(f) Baseline - With Context

### v2

@@ Abstract
As large language models (LLMs) converge towards similar capabilities, the key to advancing their performance lies in identifying and incorporating valuable new information sources. However, evaluating which text collections are worth the substantial investment required for digitization, preprocessing, and integration into LLM systems remains a significant challenge. We present a novel approach to this challenge: an automated pipeline that evaluates the potential information gain from text collections without requiring model training or fine-tuning. Our method generates multiple choice questions (MCQs) from texts and measures an LLM’s performance both with and without access to the source material. The performance gap between these conditions serves as a proxy for the collection’s information potential. We validate our approach using five strategically selected datasets: EPFL PhD manuscripts, a private collection of Venetian historical records, two sets of Wikipedia articles on related topics, and a synthetic baseline dataset. Our results demonstrate that this method effectively identifies collections containing valuable novel information, providing a practical tool for prioritizing data acquisition and integration efforts.
@@ 1 Introduction
    An automated filtering methodology for ensuring MCQ quality through complementary similarity metrics that address two key challenges: verifying question relevance and grounding in source material, while ensuring distractors are plausible yet unambiguously incorrect
@@ 2 Related Work
Information Gain. The challenge of efficiently selecting new information sources for LLMs can be viewed through the lens of optimal experiment design, a field pioneered by Fedorov et al. (1972). This framework emphasizes maximizing information gain while being strategic about resource allocation – a particularly relevant consideration given the costs associated with integrating new data into LLM systems. Information gain itself has been conceptualized across various fields: in information theory, it relates to reductions in algorithmic information content Cover et al. (1989); in machine learning, it quantifies a feature’s contribution to model performance Odhiambo Omuya et al. (2021); and in cognitive science, it represents uncertainty reduction in our experience of the world Damiano et al. (2021). While these theoretical frameworks provide valuable insights, they have not been previously applied to the specific challenge of evaluating the potential value of text collections for enhancing LLM knowledge. Our work bridges this gap by proposing a practical, MCQ-based approach that quantifies information gain by measuring an LLM’s ability to answer questions about a text collection with and without access to the source material.
@@ 3 Methods
Generating high-quality MCQs presents unique challenges, particularly due to the absence of ground truth signals typically available in human-curated educational assessments. Without explicit supervision on what constitutes a good question or appropriate distractors, we must derive reliable metrics to filter and validate the generated MCQs. This automated filtering setting raises two critical challenges: ensuring questions are both relevant and non-trivial, and guaranteeing that distractors are plausible yet unambiguously incorrect. The latter is particularly crucial when evaluating LLMs – if distractors are too simple, models might succeed through elimination rather than true knowledge; if too similar to the correct answer, they might confuse models even when provided with context. To address these challenges, we propose similarity-based filtering metrics that serve as proxy signals for question quality and difficulty, enabling automated quality control without human supervision.
@@ 3.3 Evaluation
Here, $\mathbbm{1}(\cdot)$ is an indicator function equal to 1 if the condition is true, and 0 otherwise. The denominator excludes questions that the model gets wrong both with and without context, ensuring we only measure improvement on questions where success is possible in at least one condition.
Position Bias Mitigation. To address the known issue of positional bias in LLMs Shi et al. (2024a), we evaluate each MCQ four times, rotating the correct answer through all possible positions (A, B, C, D) while randomly arranging the distractors. This effect can be seen in Figures 4, 5, and 6. This ensures that model performance reflects true knowledge rather than position-based preferences.
@@ 3.4 Data
To evaluate our pipeline’s efficacy, we conduct experiments across five strategically selected datasets representing distinct knowledge domains: (1) 177 EPFL PhD manuscripts containing specialized academic research, (2) a private collections of historical records on Venetian urban history, (3) Wikipedia articles related to EPFL manuscript topics, (4) Wikipedia articles connected to Venetian history, and (5) synthetic LLM-generated text. This dataset selection facilitates a systematic evaluation of knowledge coverage and novelty.
We hypothesize that PhD manuscripts and the Venetian historical records will yield the highest knowledge gaps, as they both contain information likely absent from training data - recent research contributions in the case of PhD manuscripts, and rare historical records in the case of the Venetian collection. The synthetic dataset, comprising model-generated content, serves as a lower-bound baseline for knowledge novelty. Wikipedia articles, being a primary source for LLM training Brown et al. (2020); Touvron et al. (2023), function as a control group representing knowledge presumably well-embedded in the models’ parameters.
Data collection. The EPFL PhD manuscripts were systematically acquired from the university’s institutional repository, which functions as the authoritative digital archive for doctoral dissertations.
The Venetian urban history collection consists of public domain books about Venice and its history. The vast majority had not been digitized previously, suggesting that they contain historical knowledge likely absent from the training data of contemporary large language models (LLMs).
The Wikipedia datasets are created by leveraging the Wikipedia API. With the help of the PhD manuscript titles and Venetian history topics, related articles are fetched and combined into similar-sized texts. Similarly, these titles and topics are used to generate the baseline dataset. To overcome the LLMs’ generation limits, in a first step, subtopics are generated surrounding the manuscript titles and Venetian history themes. Then, chunks of around 600 words are generated, using the subtopics as overarching themes, and concatenated.
@@ 4 Results
GPT-4o Wikipedia-EPFLLlama 70B BaselineLlama 70B Wikipedia-EPFLGPT-4o Wikipedia-VeniceGPT-4o EPFLLlama 70B EPFLGPT-4o Venice000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Fraction of Correct ResponsesIP ScoreBoth CorrectWith ContextNo ContextBoth Incorrect
Figure 2: Information Potential Analysis Across Datasets and Models. Stacked barplot showing correct response overlap between context-free and context-provided conditions. Overlaying the barplot is a line plot showing the Information Potential (IP) scores. Higher IP scores indicate greater novel information content, with PhD manuscripts (EPFL) showing consistently higher IP (0.211-0.229) compared to Wikipedia (0.110-0.136) and synthetic baseline (0.125). Both open and closed-source models exhibit similar patterns despite architectural differences.
Our analysis reveals three key findings: (1) similarity thresholding effectively increases question difficulty while maintaining answerable questions, (2) PhD manuscripts and Venetian records contain significantly more novel information compared to Wikipedia and synthetic datasets, and (3) larger language models show consistently lower information potential across all remaining datasets, suggesting better knowledge retention during pre-training. We present detailed evidence for each of these findings below.
@@ 4.1 Effectiveness of Similarity Thresholding
001010101020202020303030304040404050505050606060600.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111Cutoff Percentile ValuesFraction of Correct ResponsesCosineRouge-L and JaccardWC 4xNC 4xQuestions Remaining
Figure 3: Llama 3 70B Performance on EPFL Dataset Across Different Cutoff Percentages. (with Separated Thresholding and Number of MCQs for the Corresponding Percentiles). NC 4x: Evaluation with no context correct for all 4 bias mitigation evaluations of each question; WC 4x: Evaluation with context correct for all 4 bias mitigation evaluations of each question. Questions Remaining: Fraction of the original number of questions still remaining for the given cutoff threshold. Cosine: Only cosine thresholding applied. Rouge-L and Jaccard: Only Rouge-L and Jaccard thresholding applied.
Effect of Cosine Similarity Thresholding. Figure 3 displays the effect of cosine thresholding. It shows a strong decrease of around 10% between the absence of thresholding and the 50th percentile cutoff in the performance of the model with no context (NC 4x), while that of the model with context (WC 4x) remains steady. Therefore, the cosine similarity has the expected effect of increasing the MCQ difficulty when answering without context while maintaining the context-based performance.
Effect of Jaccard and ROUGE-L Thresholding. A different trend is observed when applying the Jaccard and ROUGE-L thresholding, as seen in Figure 3. Here, a slight increase in performance with context (WC 4x) of 2% can be observed between no thresholding and the 50th percentile cutoff, while performance without context (NC 4x) remains stable.
Comparative Analysis of Thresholding Methods. Our results demonstrate that cosine similarity thresholding is effective at increasing question difficulty while Jaccard and ROUGE-L slightly improve question clarity shown by the small increase in performance when the model is extended with context. Due to the separate percentile-based filtering of Jaccard and ROUGE-L scores followed by their combination, a substantial drop in the number of retained questions is observed between the 50th and 60th percentiles.
@@ 4.2 Information Potential (IP) Across Different Datasets
Limited-Access Knowledge Collections (IP: 0.229 - 0.265). Both the EPFL PhD manuscripts and Venetian historical records yield substantially higher information potential than other tested collections. We observe significant performance gaps between context-free (73.4% and 70.5%, respectively) and context-provided (98% and 97.5%) conditions, indicating considerable novel information content. These complementary results confirm our hypothesis that specialized collections with limited digital accessibility—whether recent academic research or rare historical records—represent particularly valuable information sources for LLMs, containing knowledge that is hard to retrieve or not well-represented in their training data.
Wikipedia Datasets (IP: 0.110 - 0.180). Both Wikipedia datasets demonstrate moderate information potential, with notably smaller gaps between context-free and context-provided performance (83.6% vs. 97.8% for EPFL-related Wikipedia articles and 82.8% vs. 97.2% for Venetian history Wikipedia articles). This attenuated differential suggests that a substantial portion of this information is already encoded in the models’ parameters. We hypothesize that these Wikipedia articles, or semantically equivalent content, were included in the models’ training corpora, allowing them to answer derived MCQs with relatively high accuracy even without contextual assistance. This observation corroborates our initial hypothesis and validates our methodological approach of using Wikipedia as a controlled baseline.
@@ 4.4 Qualitative Analysis of High-Value Information
Table 1: Examples of technical terminology, unique mentions, and complex relationships in EPFL PhD manuscripts. MCQs requiring context for correct model responses
@@
    Unique Mentions: The question regarding "interference errors" exemplifies how PhD manuscripts capture recent research outcomes. The detailed distinction between interference errors and reduction of intentionality errors represents novel academic insights that weren’t available during model pre-training. This type of question effectively identifies valuable new knowledge contributions from academic manuscripts. The consistency with which models fail these questions without context, despite their strong general reasoning capabilities, suggests genuine knowledge gaps rather than reasoning limitations.
@@ 5 Conclusion
    Empirical Validation: Our results validate the method’s effectiveness by demonstrating alignment with intuitive expectations across both dataset types and model scales. The information potential increases from synthetic baseline (0.125) to Wikipedia articles (0.136-0.142) to EPFL manuscripts (0.229) to Venetian historical records (0.267), while larger models (GPT-4o) consistently show lower information potential than smaller ones (Llama 70B), suggesting better knowledge retention.
@@
The successful application of our method to both PhD manuscripts and the Venetian historical collection highlights its versatility across specialized knowledge domains, reinforcing its potential as a practical tool for evaluating diverse information sources—from contemporary academic research to rare historical records. This cross-domain applicability underscores the importance of identifying areas where large language models (LLMs) struggle to retrieve information, as these “gray zones” may correspond to gaps in pretraining data. As isolating the dataset’s intrinsic contribution would require costly fine-tuning experiments, directly correlating information gain with missing pretraining data remains challenging. Nonetheless, recognizing these limitations is crucial. Just as students who are aware of their weakest subjects can better prepare for an exam—perhaps even bringing a well-structured cheat sheet if allowed—models that can identify and incorporate these gray zones as input in domain-specific tasks may achieve greater proficiency in handling specialized knowledge.
6 Limitations
Practical Considerations. Our approach, while effective, presents several methodological limitations that suggest directions for future research. First, the current pipeline’s generation-first approach, where questions are created before applying filtering criteria, could be enhanced by incorporating context-answer alignment and distractor plausibility directly into the generation process. Such an integrated approach could improve question quality while reducing computational overhead.
Second, the scope of our evaluation faces an inherent challenge: identifying datasets that are definitively outside the training distribution of large language models. While PhD manuscripts offer a reasonable proxy for novel content, the growing scale and opacity of training corpora make it increasingly difficult to verify whether any given dataset was truly unseen. This limitation underscores the urgency of developing methods to assess and prioritize valuable textual resources before they are absorbed into future training data.
Finally, our method focuses exclusively on selection rather than integration. While our results highlight where information gaps likely exist, they do not resolve how selected content might be optimally incorporated into LLMs. Emerging work, such as SKILL (Moiseev et al., 2022), proposes promising directions for efficient knowledge injection, while others (Liu et al., 2024) caution against potential interference with previously learned representations. Addressing these tensions remains an important challenge for future research.
On Human Evaluation and Pre-training Experiments. This study also omits human evaluation and pre-training-based validation. This is a deliberate design choice: our goal is to develop a lightweight, scalable method that can be used when such resource-intensive steps are impractical. Rather than replacing these approaches, our pipeline is designed as a screening mechanism. Importantly, our results exhibit strong alignment with expected knowledge gaps across datasets (e.g., low information potential for synthetic text, moderate for Wikipedia, high for novel PhD and historical records), providing empirical support for the method’s validity. Furthermore, we integrate multiple safeguards—including MCQ quality filters, position-debiased evaluation, and multi-model testing—to enhance robustness. While future work may explore downstream fine-tuning or human validation, we argue that our approach already provides reliable, cost-effective insight for strategic dataset selection.
@@ 6 Limitations
More broadly, our effort to categorize high value information sheds light on fundamental challenges in distinguishing between information that is unknown to an LLM versus information it fails to retrieve. This raises important questions about the relationship between knowledge possession and practical competence in AI systems. When an LLM consistently fails to demonstrate knowledge in a specific domain, the distinction between these cases may become less relevant from a practical perspective. This observation has broader implications for how we conceptualize and evaluate knowledge in AI systems, suggesting that performance-based metrics can also bring complementary analysis to theoretical attempts that map internal knowledge representations.
@@ A.2 Supplementary Results
In the following we present additional results for the baseline, Wikipedia-EPFL, and EPFL datasets.
@@ A.2.1 Positional Bias
ABCD001010101020202020303030304040404050505050PercentageBaselineWikipedia-EFLEPFL
@@
ABCD001010101020202020303030304040404050505050PercentageEPFL - No ContextEPFL - With ContextWikipedia-EPFL - No ContextWikipedia-EPFL - With Context
@@
ABCD001010101020202020303030304040404050505050PercentageEPFL - No ContextEPFL - With ContextWikipedia-EPFL - No ContextWikipedia-EPFL - With ContextBaseline - No ContextBaseline - With Context